OpenStudy (anonymous):

sqrt{1-3x-3=10 solve.

OpenStudy (anonymous):

yr question not clear. pls use the "Equation" button below to send your question properly

OpenStudy (anonymous):

\[\sqrt{1-3x-3}=10\]

OpenStudy (anonymous):

ok, square both sides n tell me what u get...

OpenStudy (anonymous):

???

OpenStudy (anonymous):

5 and 3

jhonyy9 (jhonyy9):

1-3x-3 =100 --- -3x-2=100 --- x=102/-3 = -34 x=-34